OpenStudy (anonymous):

PLEASE HELP Which is the equation for the line perpendicular to y = – 5/3 x +11 1/3 and containing P(–2, 3)? (1 point)

OpenStudy (anonymous):

A. Y-2 = -3/5(X-3) B. -5/3x + 4 1/3 C. -3/5x + 4 1/5 D. 3/5 x +4 1/5

OpenStudy (anonymous):

?

OpenStudy (anonymous):

The m(slope) for second line is -1/-5/3 caz we want the line perpendicular . So slope (m)is : m=-1/-5/3=1/5/3=3/5 And we want that containing P(–2, 3) . so use this : Y2-y=m(X2-x) Y2=3 X2=-2 So we have : 3-y=3/5(-2-x) -y=-6/5-3/5x y=6/5+3/5x Got it ?:) @ReaganBeard

OpenStudy (anonymous):

Yes i think

OpenStudy (anonymous):

If have any problem say ...:)

OpenStudy (anonymous):

So would it be B?

OpenStudy (anonymous):

or would it be 3/5?

OpenStudy (anonymous):

No i get it it's D

OpenStudy (anonymous):

You say : Which one of 3 and -2 are x and y ?